If X follows Bernoulli distribution Bp,p > 0.5 and V(X) 0.24 . E(X)?
Answer #1

(a)

By the property of Bernoulli Distribution:

Variance of Bernoulli Distribution is given by:

V(X) = p(1-p)

Given:
V(X) = 0.24

Thus, we get:

p (1 - p) = 0.24

i.e.,

p2 - p + 0.24 = 0

So,

1 ± V1 _ (4 × 0.24) 1 ± 0.2

Given:

p > 0.5.

So, take positive sign.

We get:

1.2 p=-=0.0

So,

Answer is:

0.6

(b)

By Propertyof Bernoulli Distribution:
E(X) = p

Thus, we get:
E(X) = p = 0.6

So,

Answer is:

0.6